Solve system by elimination:
3x-2y=-1
3x-4y=9

Answer by checkley77(12844)

3x-2y=-1
-(3x-4y=9) subtract
-----------------
2y=-10
y=-10/2
y=-5 ans.
3x-2*-5=-1
3x+10=-1
3x=-1-10
3x=-11
x=-11/3 ans.
Proof:
3*-11/3-4*-5=9
-11+20=9
9=9